A shopkeeper sold 100 fruits (3 types). Apple was sold for 10 cents each,Banana 1 Euro each and Orange sold for 3 Euro each. The total amount collected for the 100 fruits was 100 Euros. How many fruits of each type were sold? thanks
Answer by ankor@dixie-net.com(22740)

A shopkeeper sold 100 fruits (3 types). Apple was sold for 10 cents each,Banana 1 Euro each and Orange sold for 3 Euro each. The total amount collected for the 100 fruits was 100 Euros. How many fruits of each type were sold?
Let a = no. of apples, b = no. of banana, r = no. of oranges
:
Since Banana's & Organges are in Euros, the number of apples sold at .10 have to
be multiples of 10 in order for it to add up to 100 Euros.
I tried 30 apples without success, however with 40 apples:
:
The number of fruit equation
40 + b + r = 100
b + r = 100 - 40
b + r = 60
:
And the Euro amt equation
.10(40) + 1b + 3r = 100
4 + 1b + 3r = 100
b + 3r = 100 - 4
b + 3r = 96
:
Gives us two, 2 unknown equations, we can solve by elimination:
b + 3r = 96
b + r = 60
------------- subtract
0 + 2r = 36
r = 36/2
r = 18 oranges
:
Substitute 40 for a and 18 for r in the fruit quantity equaiton:
40 + b + 18 = 100
b = 100 - 58
b = 42 bananas
:
Check our solutions in the euro equation:
.10(40) + 1(42) + 3(18) =
4 + 42 + 54 = 100
:
Solution: 40 apples, 42 bananas and 18 oranges